# Importance-Aware Robust Semantic Transmission for LEO Satellite-Ground Communication

### v1

@@ Abstract
Satellite-ground semantic communication is anticipated to serve a critical role in the forthcoming 6G era. Nonetheless, task-oriented data transmission in such systems remains a formidable challenge, primarily due to the dynamic nature of signal-to-noise ratio (SNR) fluctuations and the stringent bandwidth limitations inherent to low Earth orbit (LEO) satellite channels. In response to these constraints, we propose an importance-aware robust semantic transmission (IRST) framework, specifically designed for scenarios characterized by bandwidth scarcity and channel variability. The IRST scheme begins by applying a segmentation model enhancement algorithm to improve the granularity and accuracy of semantic segmentation. Subsequently, a task-driven semantic selection method is employed to prioritize the transmission of semantically vital content based on real-time channel state information. Furthermore, the framework incorporates a stack-based, SNR-aware channel codec capable of executing adaptive channel coding in alignment with SNR variations. Comparative evaluations across diverse operating conditions demonstrate the superior performance and resilience of the IRST model relative to existing benchmarks.
Index Terms:
Semantic communication, satellite-ground communication, image transmission, semantic selective, adaptive coding.
@@ I Introduction
As the global demand for high-speed, low-latency, wide-coverage communications continues to grow, satellite-ground communication will play an indispensable role in future 6G era [1, 2]. Compared to traditional terrestrial communication networks, satellite-ground communication has a wider coverage capacity, especially in remote mountain areas, oceans, and polar regions where the ground infrastructure is difficult to deploy, greatly improving the universality and accessibility of communication [3]. Moreover, satellite-ground communication, particularly low-Earth orbit (LEO) satellite communication, offers several compelling advantages, including reduced communication latency, lower path loss, enhanced system capacity, and greater operational flexibility [4, 5, 6]. Therefore, satellite-ground communication plays an important role in emerging application scenarios, including the global Internet of Things (IoT), disaster response, and telemedicine [7].
@@
Driven by these advantages, researchers have employed SemCom in satellite-ground systems. As mentioned above, the dynamic change of the satellite-ground channel affects the stability of communication and presents the challenge of deterioration of link reliability. To address the issue, Chen et al. [15] propose a method combining SemCom with orthogonal time-frequency space (OTFS) modulation to effectively alleviate the severe Doppler effect and improve transmission efficiency. Furthermore, Jiang et al. [17] use multirandom noise and multichannel mulation to improve the robustness of the transmission data during the training process, so that a single training model can be adapted to remote sensing applications with different signal-to-noise ratios (SNR) and harsh environments. Yuan et al. [18] propose a CSI-guided ModNet channel that aligns image semantics with transmission conditions, optimizing the trade-off between reconstruction quality and network utility. Moreover, Jiang et al. [19] investigate denoising and restoration strategies for satellite-based SemCom networks. Among these, FM-based segmentation and reconstruction techniques demonstrate robust performance in preserving semantic features under high-noise conditions. In addition, advances have been made in optimizing the selection of gateway hop relays for low-SNR ground users, coupled with the integration of gateway-based denoising mechanisms [20].
@@
Furthermore, the bandwidth resource of the satellite-ground communication system is limited, reducing the utilization efficiency. To address this issue, semantic importance is a feasible method. Sun et al. [21] calculate semantic importance by analyzing the gradient of task perception results, and incorporate semantic weighting into the loss function design. To overcome related limitations, various semantic awareness strategies have emerged, emphasizing data utility and relevance across interconnected devices [22]. Approaches include the FAST framework for feature evaluation [23], entropy-based models for latent semantic discrimination [24], and gradient-based methods for semantic concept importance [25, 26]. Features exceeding a predefined weight threshold are preserved, while others are discarded.
@@
However, the aforementioned methods are not directly applicable to the satellite-ground scene. First, semantic evaluation models trained in terrestrial environments often struggle to generalize to the unique noise and interference characteristics of satellite-ground channels, leading to biased semantic weight assessments and degraded transmission efficiency [19]. Second, many semantic awareness strategies (such as gradient-based analysis and entropy modeling) are computationally intensive [22], while satellites typically operate under constrained processing capabilities and energy budgets, limiting their ability to perform real-time reasoning or update model parameters [27].
@@
To address the aforementioned challenges, we propose an importance-aware robust semantic transmission (IRST) model that enables content-aware transmission under dynamic SNR conditions. The proposed framework leverages semantic-level selection mechanisms in conjunction with enhanced channel coding strategies to adaptively prioritize and transmit information that is most relevant to the communication task. Table I compares the proposed scheme with existing representative SemCom schemes. The main contributions of this paper are as follows:
@@ II-C1 Semantic Segmentation Module
An RGB image source $\mathbf{s}$ is input to the semantic segmentation model. The model performs semantic segmentation and generates the corresponding output $\mathbf{s_{ss}}$. This process is modeled as
@@ II-C2 Semantic Selection Module
where $x$ is the image that needs to be transmitted, and $\gamma$ and q{q} are the current SNR and the task requirements, respectively.
@@ II-C3 Data Transmission Module
where $\mathbf{x_{se}}$ is the result of the semantic encoder. $\mathbf{\theta_{se}}$ and $\mathbf{\theta_{ce}}$ are the model parameters of the semantic encoder and the channel encoder, respectively. Beyond the encoder, $\mathbf{y}$ will be sent to the satellite-ground channel for transmission. Due to the existence of channel fading and noise, the signal $\mathbf{y}$ will change to $\mathbf{\tilde{y}}$ after passing through the channel. The fading process is modeled as
@@ III-A1 Encoder
The encoder uses the first 13 convolutional layers of VGG16 for feature extraction. This step uses standard 2D convolution, which can be expressed as
$\begin{array}[]{l}Y(i,j)=\sum\limits_{m=0}^{M-1}{\sum\limits_{n=0}^{N-1}{X(i+m,j+n)\cdot W(m,n)+b}},\end{array}$ (9)
where $X(i,j)$ is the input feature map, $W(m,n)$ is the convolution kernel, $b$ is the bias term, and $Y(i,j)$ is the output feature map. Then, the resolution is reduced by a 222*2 maximum pooling to reduce the amount of calculation. This step can be expressed as
$\begin{array}[]{l}Y(i,j)=max\{X(2i,2j),X(2i+1,2j),X(2i,2j+1),\\ \hskip 40.0ptX(2i+1,2j+1)\}.\end{array}$ (10)
@@ III-A2 Decoder
Here, $X(i^{\prime},j^{\prime})$ represents the eigenvalue obtained from the encoder, and $X^{\prime}(i,j)$ represents the upsampling result after reconstruction in the decoder. Compared with the traditional deconvolution or bilinear interpolation methods, this mechanism has stronger geometric structure retention ability and can complete the recovery of spatial resolution without introducing redundant information [39].
@@ IV-A Transmission Effect Evaluation Scheme
IV-B Semantic Selection Process
@@
Firstly, the semantic segmentation model is used to process the input image, identify different categories, and apply color mapping. Then, according to the different tasks, the target area of the task is screened by mask operation to generate an adjusted binary image. On this basis, the transmission effect evaluation model is used to evaluate the recovery of the image after channel transmission in the current noise environment. According to the evaluation results, different mean fuzzy filtering schemes are used to adaptively adjust the non-main part of the image to enhance the visual effect and improve the accuracy of the analysis. Finally, in combination with mask correction technology, an optimized image is generated to prepare for subsequent data transmission. Figure 3 shows the effect.
@@ V-A Semantic Codec Architecture
In the encoder part, the role of the encoder is to extract multi-scale features, which is mainly composed of Swin Transformer blocks and patch merging. The Swin Transformer blocks part adopts the window self-attention (W-MSA) mechanism and the shift window self-attention (SW-MSA) mechanism [38]. W-MSA is used to calculate self-attention in a local window to improve computational efficiency. SW-MSA realizes the interaction of information across windows through the window shift mechanism. The patch merging part adopts the gradual merging of adjacent patches to improve the feature expression ability by reducing the spatial dimension. At the same time, this part adopts a hierarchical structure, so that the model can learn features of different scales.
@@ VI Simulation Results and Analysis
VI-A Dataset and Simulation Parameters

### v2

@@ Abstract
Satellite-ground semantic communication is anticipated to serve a critical role in the forthcoming sixth-generation (6G) mobile networks. Nonetheless, task-oriented data transmission in such systems remains a formidable challenge, primarily due to the dynamic nature of Signal-to-Noise Ratio (SNR) fluctuations and the stringent bandwidth limitations inherent to Low Earth Orbit (LEO) satellite channels. In response to these constraints, we propose an Importance-Aware Robust Semantic Transmission (IRST) framework, specifically designed for scenarios characterized by bandwidth scarcity and channel variability. The IRST scheme begins by applying a segmentation model enhancement algorithm to improve the granularity and accuracy of semantic segmentation. Subsequently, a task-driven semantic selection method is employed to prioritize the transmission of semantically vital content based on real-time Channel State Information (CSI). Furthermore, the framework incorporates a stack-based, SNR-aware channel codec capable of executing adaptive channel coding in alignment with SNR variations. Comparative evaluations across diverse operating conditions demonstrate the superior performance and resilience of the IRST model relative to existing benchmarks. The code is available at https://github.com/lightwindy-ch/IRST.git.
@@ I Introduction
As the global demand for high-speed, low-latency, wide-coverage communications continues to grow, satellite-ground communication will play an indispensable role in sixth-generation (6G) mobile networks [1, 2]. Compared to terrestrial networks, satellite-ground communication offers broader coverage, particularly in regions with limited infrastructure such as mountains, oceans, and polar areas, thereby enhancing communication accessibility [3]. Moreover, satellite-ground communication, particularly Low Earth Orbit (LEO) satellite communication, offers several compelling advantages, including reduced communication latency, lower path loss, enhanced system capacity, and greater operational flexibility [4, 5, 6]. Therefore, satellite-ground communication plays an important role in emerging application scenarios, including the global Internet of Things (IoT), disaster response, and telemedicine [7].
@@
Driven by these advantages, researchers have employed SemCom in satellite-ground systems. As mentioned above, the dynamic change of the satellite-ground channel affects the stability of communication and presents the challenge of deterioration of link reliability. To address the issue, Chen et al. [15] propose a method combining SemCom with Orthogonal Time-Frequency Space (OTFS) modulation to effectively alleviate the severe Doppler effect and improve transmission efficiency. Furthermore, Jiang et al. [18] use multirandom noise and multichannel mulation to improve the robustness of the transmission data during the training process, so that a single training model can be adapted to remote sensing applications with different Signal-to-Noise Ratio (SNR) and harsh environments.Yuan et al. [19] propose a ModNet channel guided by Channel State Information (CSI) that aligns the semantics of the image with the transmission conditions, optimizing the trade-off between the quality of the reconstruction and the utility of the network. Moreover, Jiang et al. [20] investigate denoising and restoration strategies for satellite-based SemCom networks. Among these, Foundation Model (FM)-based segmentation and reconstruction techniques demonstrate robust performance in preserving semantic features under high-noise conditions. In addition, advances have been made in optimizing the selection of gateway hop relays for low-SNR ground users, coupled with the integration of gateway-based denoising mechanisms [21].
@@
Furthermore, the bandwidth resource of the satellite-ground communication system is limited, reducing the utilization efficiency. To address this issue, semantic importance is a feasible method. Sun et al. [22] calculate semantic importance by analyzing the gradient of task perception results, and incorporate semantic weighting into the loss function design. To overcome related limitations, various semantic awareness strategies have emerged, emphasizing data utility and relevance across interconnected devices [23]. Approaches include the Feature Allocation for Semantic Transmission (FAST) framework for feature evaluation [24], entropy-based models for latent semantic discrimination [25], and gradient-based methods for semantic concept importance [26, 27]. Features exceeding a predefined weight threshold are preserved, while others are discarded.
@@
However, the aforementioned methods are not directly applicable to the satellite-ground scenario. First, semantic evaluation models trained in terrestrial environments often struggle to generalize to the unique noise and interference characteristics of satellite-ground channels, leading to biased semantic weight assessments and degraded transmission efficiency [20]. Second, many semantic awareness strategies (such as gradient-based analysis and entropy modeling) are computationally intensive [28], while satellites typically operate under constrained processing capabilities and energy budgets, limiting their ability to perform real-time reasoning or update model parameters [29].
Therefore, although SemCom has made significant progress in terrestrial networks, its application in satellite-ground communication scenarios still faces many challenges. Especially in terms of dynamic noise reduction mechanism, satellite-ground channel adaptability, and multi-dataset generalization ability, the existing methods have obvious limitations. Most existing methods rely on fixed semantic compression and recovery structures, lacking adaptability to real-time channel variations. Secondly, in the satellite-ground link, channel noise may lead to misjudgment of semantic features, thereby incorrectly discarding key semantic information and affecting task completion. Meanwhile, inconsistent definitions of semantic importance across datasets hinder the development of a unified selection mechanism. Finally, most methods rely on offline training and lack a lightweight online update mechanism, which cannot adapt to task changes or environmental disturbances.
@@
To address the aforementioned challenges, we propose an Importance-Aware Robust Semantic Transmission (IRST) model that enables content-aware transmission under dynamic SNR conditions. The framework integrates three modules: semantic segmentation, semantic selection, and data transmission, achieving end-to-end optimization from semantic extraction to channel adaptation. This framework not only improves semantic recovery capabilities in low SNR environments, but also significantly reduces redundant data transmission, and has stronger task adaptability and cross-scenario generalization capabilities. Meanwhile, in contrast to existing semantic perception methods that are computationally intensive, the proposed IRST framework achieves substantial complexity reduction through a modular architecture and dynamic adaptation strategies. Specifically, the semantic selection module employs a lightweight evaluation function to identify task-relevant regions, thereby circumventing the need for gradient-based optimization or entropy-based modeling. Furthermore, the SNR-adaptive channel codec incorporates a hierarchical activation mechanism that selectively engages deeper network layers only under low-SNR conditions, effectively conserving computational resources in high-SNR scenarios. Finally, the framework supports end-to-end inference without requiring frequent model updates or retraining, rendering it well-suited for deployment in resource-constrained satellite platforms.
Table I compares the proposed scheme with existing representative SemCom schemes. The main contributions of this paper are as follows.
@@ II-C1 Semantic Segmentation Module
An image source of Red, Green, Blue (RGB) $\mathbf{s}$ is input to the semantic segmentation model. The model performs semantic segmentation and generates the corresponding output $\mathbf{s_{ss}}$. This process is modeled as
@@ II-C2 Semantic Selection Module
where $x$ is the image that needs to be transmitted, and $\gamma$ and q{q} are the current SNR in wireless channel and the task requirements, respectively.
@@ II-C3 Data Transmission Module
where $\mathbf{x_{se}}$ is the result of the semantic encoder. $\mathbf{\theta_{se}}$ and $\mathbf{\theta_{ce}}$ are the model parameters of the semantic encoder and the channel encoder, respectively. $\mathbf{y}$ is the signal after passing through the channel encoder. Beyond the encoder, $\mathbf{y}$ will be sent to the satellite-ground channel for transmission. Due to the existence of channel fading and noise, the signal $\mathbf{y}$ will change to $\mathbf{\tilde{y}}$ after passing through the channel. The fading process is modeled as
@@
In summary, the IRST framework emphasizes the coordination and feedback mechanism between modules. The semantic segmentation module uses the SME algorithm to enhance boundary recognition, providing more accurate semantic labels for subsequent semantic selection. The semantic selection module dynamically adjusts the importance distribution of image content based on task requirements and channel conditions. The data transmission module selects the optimal coding path based on real-time channel quality, achieving a balance between computational resources and transmission efficiency. These three components work together to build a complete communication link with semantic perception, task-driven and channel adaptive ability.
@@ III-A1 Encoder
The encoder uses the first 13 convolutional layers of Visual Geometry Group 16 (VGG16) for feature extraction. This step uses standard 2D convolution, which can be expressed as
$\begin{array}[]{l}T(i,j)=\sum\limits_{m=0}^{M-1}{\sum\limits_{n=0}^{N-1}{S(i+m,j+n)\cdot W(m,n)+b}},\end{array}$ (9)
where $S(i,j)$ is the input feature map, $W(m,n)$ is the convolution kernel, $b$ is the bias term, and $T(i,j)$ is the output feature map. Then, the resolution is reduced by a 222*2 maximum pooling to reduce the amount of calculation. This step can be expressed as
$T(i,j)=\max\big\{$ $S(2i,2j),\,S(2i+1,2j),\,S(2i,2j+1),$ (10)
$\hskip 40.0ptS(2i+1,2j+1)\big\}$
@@ III-A2 Decoder
Here, $S(i^{\prime},j^{\prime})$ represents the eigenvalue obtained from the encoder, and $S^{\prime}(i,j)$ represents the upsampling result after reconstruction in the decoder. Compared with the traditional deconvolution or bilinear interpolation methods, this mechanism has stronger geometric structure retention ability and can complete the recovery of spatial resolution without introducing redundant information [41].
@@ III-B Segmentation Enhancement Scheme
Then, let us explain the SME algorithm from a theoretical point of view. Let $z(i,j)$ denote the pixel at position $(i,j)$, and $\mathcal{N}(i,j)$ denote its $3\times 3$ neighborhood excluding the center. Define the majority function as
$\begin{array}[]{l}M(i,j)=\arg\max_{c\in\mathcal{C}}\;\text{count}_{\mathcal{N}(i,j)}(c),\end{array}$ (13)
where $\mathcal{C}$ is the set of color values. The SME update rule is modeled as
$\begin{array}[]{l}z^{\prime}(i,j)=\begin{cases}M(i,j),&\text{if }\max_{c}\text{count}_{\mathcal{N}(i,j)}(c)\geq\tau,\\ z(i,j),&\text{otherwise},\end{cases}\end{array}$ (14)
with threshold $\tau=7$.
This mechanism can be interpreted as a local denoising operator that minimizes the probability of edge misclassification. Specifically, if the probability of noise corruption per pixel is $p$, the probability that a majority of neighbors are corrupted is bounded by
$\begin{array}[]{l}{P_{error}}\leq\sum\limits_{k=\tau}^{8}{\left(\begin{array}[]{c}8\\ k\end{array}\right)}{p^{k}}{(1-p)^{8-k}}.\end{array}$ (15)
with threshold $\tau=7$.
Thus, SME significantly reduces edge error when $p<0.3$, ensuring the accuracy in semantic segmentation.
@@ IV-A Transmission Effect Evaluation Scheme
IV-B Theoretical Modeling of Semantic Selection
To enhance the theoretical rigor of the semantic selection module, we formalize the selection process as a task-driven optimization problem under channel constraints. Let the original image be denoted by $\mathbf{s}\in\mathbb{R}^{H\times W\times C}$, and the semantic segmentation output by $\mathbf{s_{ss}}\in\mathbb{R}^{H\times W\times K}$, where $K$ is the number of semantic categories. The task requirement is represented by a weight vector $q\in\mathbb{R}^{K}$, and the current channel condition (e.g. SNR) is denoted by $\gamma\in\mathbb{R}$.
We define the semantic relevance score $\mathcal{I}(i,j)$ for each pixel location $(i,j)$ as:
$\mathcal{I}(i,j)=\sum_{k=1}^{K}\mathbf{s_{ss}}(i,j,k)\cdot q_{k},$ (21)
where $\mathbf{s_{ss}}(i,j,k)$ is the probability that pixel $(i,j)$ belongs to category $k$, and $q_{k}$ reflects the task-specific importance of category $k$.
Given the limited transmission budget $B(\gamma)$, which is a function of the channel condition $\gamma$, the selection of image regions $x\subset s$ must satisfy:
$\sum_{(i,j)\in x}\mathcal{C}(i,j)\leq B(\gamma),$ (22)
where $\mathcal{C}(i,j)$ denotes the transmission cost of pixel $(i,j)$, approximated by its encoding complexity or bit-length.
The semantic selection task is thus cast as the following constrained optimization problem:
$\max_{x\subset s}\sum_{(i,j)\in x}\mathcal{I}(i,j)\quad\text{s.t.}\quad\sum_{(i,j)\in x}\mathcal{C}(i,j)\leq B(\gamma).$ (23)
To reduce computational overhead and enable real-time deployment on resource-constrained satellite platforms, we propose a lightweight scoring function $\mathcal{S}(i,j)$ that incorporates channel awareness:
$\mathcal{S}(i,j)=\mathcal{I}(i,j)\cdot\phi(\gamma),$ (24)
where $\phi(\gamma)$ is a channel-adaptive modulation function. For instance, a threshold-based formulation may be adopted:
$\phi(\gamma)=\begin{cases}1,&\gamma\geq\gamma_{th},\\ \alpha,&\gamma<\gamma_{th},\end{cases}$ (25)
with $\alpha<1$ serving to suppress low-importance regions under poor channel conditions.
Based on $\mathcal{S}(i,j)$, the image is selectively compressed via a spatial filtering function $\mathcal{B}(s;\mathcal{S},\tau)$, where $\tau$ is a tunable threshold. Pixels with $\mathcal{S}(i,j)<\tau$ are blurred or downsampled, while high-score regions are preserved in full fidelity.
This theoretical framework enables the semantic selection module to dynamically balance task relevance and transmission efficiency, while maintaining computational tractability. It also provides a principled basis for future extensions such as multi-task prioritization and reinforcement-based selection strategies.
IV-C Semantic Selection Process
@@
Firstly, the semantic segmentation model is used to process the input image, identify different categories, and apply color mapping. Then, according to the different tasks, the target area of the task is screened by mask operation to generate an adjusted binary image. On this basis, the transmission effect evaluation model is used to evaluate the recovery of the image after channel transmission in the current noise environment. In this scheme, we use the pilot-based minimum mean square error (MMSE) channel estimation method to recover the channel response, calculate the instantaneous SNR at the receiver, and return the value to the satellite.
According to the evaluation results, different mean fuzzy filtering schemes are used to adaptively adjust the non-main part of the image to enhance the visual effect and improve the accuracy of the analysis. Finally, in combination with mask correction technology, an optimized image is generated to prepare for subsequent data transmission. Figure 3 shows the effect. In this way, according to the different transmission tasks, the proposed solution can reduce 10%-80% of the transmission content and greatly save bandwidth.
@@ V-A Semantic Codec Architecture
In the encoder part, the role of the encoder is to extract multi-scale features, which is mainly composed of Swin Transformer blocks and patch merging. The Swin Transformer blocks part adopts the Window-based Multi-head Self-Attention (W-MSA) mechanism and the Shifted Windows Multi-head Self-Attention (SW-MSA) mechanism [40]. W-MSA is used to calculate self-attention in a local window to improve computational efficiency. SW-MSA realizes the interaction of information across windows through the window shift mechanism. The patch merging part adopts the gradual merging of adjacent patches to improve the feature expression ability by reducing the spatial dimension. At the same time, this part adopts a hierarchical structure, so that the model can learn features of different scales.
@@ V-B Dynamic SNR Channel Codec Architecture
To formalize the adaptive transmission process under dynamic SNR conditions, we define the end-to-end transmission pipeline as a composition of semantic encoding, channel encoding, channel fading, and decoding operations.
The encoder is segmented into $L$ layers, each with activation condition $\gamma_{l}$ based on the current SNR $\gamma$. The activation function is defined as:
$\delta_{l}(\gamma)=\begin{cases}1,&\gamma<\gamma_{l},\\ 0,&\gamma\geq\gamma_{l},\end{cases}$ (26)
where $\delta_{l}(\gamma)$ indicates whether layer $l$ is activated under SNR $\gamma$.
The overall quality of transmission is evaluated using a semantic fidelity metric $\mathcal{F}(\mathbf{\tilde{x}},\mathbf{x})$, which may include perceptual similarity and task-specific precision. The transmission optimization objective is as follows:
$\max_{\theta_{se},\theta_{ce}}\mathbb{E}_{\gamma,h,n}\left[\mathcal{F}(\mathbf{\tilde{x}},\mathbf{x})\right]\quad\text{s.t.}\quad\mathcal{C}(y)\leq B(\gamma),$ (27)
where $\mathcal{C}(y)$ denotes the transmission cost and $B(\gamma)$ is the bandwidth budget under SNR $\gamma$.
@@ VI Simulation Results and Analysis
VI-A Theoretical Complexity Analysis
To rigorously assess the computational efficiency of the proposed IRST framework, we present a theoretical analysis of the time complexity associated with its three principal modules: the semantic segmentation module, the semantic selection module, and the data transmission module. This analysis elucidates the computational demands of each component under varying image resolutions and channel conditions, thereby establishing a theoretical foundation for subsequent runtime benchmarking.
VI-A1 Semantic Segmentation Module
The semantic segmentation module adopts the SegNet architecture, and its encoder part is based on the first 13 layers of VGG16 convolutional network. The time complexity of each layer convolution operation is $O(H\cdot W\cdot{K^{2}}\cdot{C_{in}}\cdot{C_{out}})$. Among them, $H$,$W$ are the size of the feature map, $K$ is the size of the convolution kernel, and $C_{in}$, $C_{out}$ are the size of the convolution kernel. The decoder uses upsampling based on pooling index, and the complexity is $O(H\cdot W)$. The classification layer uses Softmax to classify each pixel with a complexity of $O(H\cdot W\cdot C)$, where $C$ is the number of categories. Therefore, the overall complexity of the semantic segmentation module is approximately $O(H\cdot W\cdot C^{2})$.
VI-A2 Semantic Selection Module
The module evaluates and blurs the image region according to the task requirements and the current SNR. Due to the use of a lightweight pixel neighborhood statistical method, the complexity mainly comes from the local window operation for each pixel (such as 3 × 3 neighborhood), so the overall complexity is $O(H\cdot W)$.
VI-A3 Data Transmission Module
The module includes Swin Transformer semantic encoder and SNR segmented channel encoder. The window attention mechanism of Swin Transformer reduces the complexity from $O(N^{2}\cdot d)$ of traditional Transformer to $O(N\cdot d\cdot w^{2})$. Among them, $N=H\cdot W$, $d$ is the feature dimension, and $w$ is the window size. The channel encoder dynamically activates sub-networks of different depths according to SNR conditions. In the worst case, the complexity is $O(H\cdot W\cdot D)$, where $D$ is the maximum coding depth. Under high SNR conditions, some paths can be skipped to further reduce the computational load.
In summary, the IRST framework demonstrates effective control over computational complexity through modular architecture and conditional activation strategies. This design ensures robust semantic representation while offering adaptability to deployment constraints and runtime efficiency.
VI-B Dataset and Simulation Parameters